# A kinematic origin of the white dwarfs in the Solar neighborhood

### v1

@@ Abstract
Context. White dwarfs are considered to be feasible cosmochronometers. Thanks to the recent space-borne mission Gaia, there is a nearly complete sample up to about 100 parsecs from the Sun. However, the Galaxy, as a dynamic system, implies that these objects may have very diverse origins. It is therefore of paramount importance to characterize the origins of white dwarfs from the different Galactic structure components found in our Solar neighborhood.
Aims. We aim to compute the Galactic orbits for white dwarfs of the thin and thick disk, and halo components observed in our Solar neighborhood. Once these orbits are determined, we will analyze the most probable regions of the Galaxy where they could have formed, the distribution of their orbital parameters, as well as the observational biases introduced when constructing the local sample.
Methods. We used a detailed Galactic orbit integration package, in conjunction with a detailed population synthesis code specifically designed to replicate the different Galactic components of the white dwarf population. Synthetic stars are generated based on the current observational sample, and subsequently, their orbital integration allows for the reconstruction of the population past history.
Results. Our kinematic analysis of the white dwarf population reveals the ephemeral nature of the concept of Solar neighborhood, as the majority of thin disk, thick disk, and halo white dwarfs will have left our 100 pc neighborhood in approximately 3.30 Myr, 1.05 Myr, and 0.6 Myr, respectively. Moreover, the spatial distribution of the integrated thin disk orbits suggests that 68% of these stars were formed at less than 1 kpc from the Sun, while most of thick disk suffered from radial disk migration. Halo members are those stars that typically belong to the “inner halo”, being their orbits mostly planar and not extending beyond –25 kpc. Despite the observational bias, which mostly affects the oldest stars in the thick disk and halo, we show that the wider distribution of orbital parameters is well represented by the observed sample.
Conclusions. The Solar neighborhood is a transitory concept, whose current population of white dwarfs originates from larger regions of the Galaxy. This fact must be taken into account when analyzing the overall properties of such a population, such as its age distribution, metallicities or formation history. Even so, the kinematic properties observed by recent missions such as Gaia are representative of the total population up to a distance of approximately 500 pc.
@@
The high precision astrometry provided by the European Space Agency Gaia mission (Gaia Collaboration et al., 2016) has opened a wide parameter space to Galactic studies, enabling the mapping of stellar kinematics over a large volume of the Milky Way extending up to 6 kpc towards the inner and outer Galactic radial directions (Gaia Collaboration et al., 2018). The complementary information that is added by precise measurements of chemical abundances for ever-growing samples of stars is contributing to shape the correlations between kinematics and stellar chemistry that are key to unravel the past history of our Galaxy (Gaia Collaboration et al., 2023). The improved astrometry delivered by the Gaia Early Data Release 3 (EDR3; Gaia Collaboration et al., 2021) has confirmed the interplay between the kinematics of local structures, such as the moving groups (Antoja et al., 2008), and that of more distant ones (Ramos et al., 2018; Khanna et al., 2019) in the Galactic anti-center direction and away from the mid-plane of the Galactic disk, thus implying a connection with the recent passing of the Sagittarius dwarf galaxy (Ruiz-Lara et al., 2020) and interactions with the Large Magellanic Cloud (Antoja et al., 2018; Laporte et al., 2018).
@@
Pioneering works focused on the kinematic analyses of white dwarf samples in the pre-Gaia era (e.g., Silvestri et al., 2001; Pauli et al., 2006; Wegg & Phinney, 2012; Sion et al., 2014; Anguiano et al., 2017). However, more recently, by means of accurate Gaia parallaxes and proper motions, it has been possible to associate white dwarfs to known stellar streams (Torres et al., 2019a), to measure the reflex Solar motion (Rowell & Kilic, 2019), to assign their membership to Galactic components within 100 pc (Torres et al., 2019b), and to accurately measure their age – velocity dispersion relation (Cheng et al., 2019; Raddi et al., 2022).
In this paper, we model the kinematics of white dwarfs with a synthetic population, by integrating their Galactic orbits in a static potential and using the present-day boundary conditions that are obtained from the Gaia data. While a static potential is far from being a realistic approximation, the kinematics of local stars carries the imprint of resonances and radial mixing due to the Galactic structure and dynamics (Sellwood & Binney, 2002; Dehnen, 2000; Hunt et al., 2019), thus enabling a statistical study of the distributions that characterize their orbital parameters. The crossing-time of white dwarfs that are passing through the Solar neighborhood is numerically quantified, as it is known to introduce a bias, although small, in the average properties of classes of local stars (Mayor et al., 1977). Moreover, we analyze the spatial distributions that are obtained by integrating for several orbits around the Galaxy, as a statistical way for linking the observed white dwarf populations to the birth sites of their main-sequence progenitors.
In the following sections, we introduce the properties of the observed and simulated samples (Sec. 2 and 3, respectively), describe the orbit integration framework (Sec. 4), and present a detailed analysis of the results (Sec. 5). In the last section, we conclude with a summary and the expectations for future developments.
@@ 2 Observed sample
The analysis of the white dwarf population provided by Gaia reveals that the largest sample that can be considered as effectively complete is the 100 pc volume-limited sample (Jiménez-Esteban et al., 2018). Thus, we started from the sample analyzed in Torres et al. (2019b), in which a Random Forest algorithm was employed to classify the white dwarf population within 100 pc of the Sun into its Galactic components. The sample was selected by applying color-magnitude cuts to the Hertzsprung–Russell diagram and adding specific selection criteria, such as, photometric and astrometric relative errors under 10 per cent (see section 4 from Torres et al., 2019b, for a detailed description). Then, we cross-matched it with the Gaia EDR3, providing updated astrometry for the same sample.
@@
It is worth noting here that, in order to eliminate possible contaminants in high-crowded areas, the observed sample analyzed in Torres et al. (2019b) did not include objects with proper motions corresponding to projected tangential velocities below km s-1. This fact should be taken into account in our analysis because it may introduce a slight bias for low speed objects that will be absent in the observed sample.
@@
The observed sample was thus classified using a Random Forest algorithm applied to a 8D space (equatorial coordinates, parallax, proper motion components, and photometric magnitudes). The updated sample, with Gaia EDR3 astrometry, delivers 11,568 thin disk, 1317 thick disk, and 89 halo white dwarf candidates, implying a reduction of 6 % for each of the populations with respect to the number of candidates counted by Torres et al. (2019b). Such a small difference is due to stars now being outside the 100 pc volume. We note that we did not search for new 100-pc candidates in EDR3.Although radial velocity measurement are absent in the Torres et al. (2019b) sample, it can be considered the largest, nearly complete, sample of white dwarfs, thus justifying a reliable analysis of the kinematics of the Galactic components of this population.
@@
Since most of the known white dwarfs do not have accurate radial velocity measurements, with very few of them having suitable high-resolution observations (Napiwotzki et al., 2020), a radial velocity = 0 km s-1has been usually adopted in kinematic studies of this population. Alternative methods to overcome this issue have been proposed (Dehnen & Binney, 1998); however, they are not exempt from certain limitations (McMillan & Binney, 2009). In order to see what is the effect of considering the radial velocity zero, Torres et al. (2019b), analyzed the velocity component distributions in the Toomre diagram for the synthetic sample of the different Galactic components by subtracting the . Their analysis showed an average bias of 10 km s-1 and up to 74 km s-1 for the average component of the disk and halo stars, respectively, when the km s-1. The bias resulting by the assumption of zero radial velocity for white dwarfs is estimated to cause a reduction by a factor of one-third in the three velocity components of the heliocentric Cartesian reference system (e.g. Rowell & Kilic, 2019). Although it is beyond the scope of the present work to retrieve the full 3D velocity space of the white dwarf population, we address this issue in Section 5.2.2, in which we attempt to correct for this bias using our synthetic model.
3 The synthetic white dwarf population
@@
Halo |
@@
In addition to the observational sample just described, we produced a synthetic sample of white dwarfs to address potential observational biases or deficiencies that the observational sample may have. We took advantage of the consolidated population synthesis code based on Monte Carlo techniques described in Torres et al. (2019b), which has been widely used over the last few decades in the study of the disk, halo, Galactic bulge, open clusters, and globular clusters (e.g., Garcia-Berro et al., 1998; Torres et al., 2002; García-Berro et al., 2010; Cojocaru et al., 2015; Torres et al., 2015; Torres & García-Berro, 2016; Torres et al., 2018)
@@
The Monte Carlo algorithm provides the initial position of each star within a 100 pc radius around the Sun and its age. In what follows we provide some of the basic ingredients used in our population synthesis code. According to a certain star formation rate (SFR), a probabilistic distribution determines the the ages of stars belonging to each Galactic component. We assume a uniform SFR for the thin disk that lasts , and Gaussian distributions for the thick disk (, 1 Gyr), and the halo (, 0.5 Gyr), with 9.2, 10.0 and 12.5 Gyr, respectively (Torres et al., 2019b).
@@
The Local Standard of Rest (LSR) is the chosen reference frame to describe the positions and velocities of the white dwarfs. It is instantaneously located on the Sun at kpc from the Galactic center, performing a circular orbit rotating at km s-1, which is nearly the average rotation velocity of stars in the Solar neighborhood (Bovy et al., 2012). We place the Sun at pc above the Galactic plane (e.g., Jurić et al., 2008). Nevertheless, the actual values for the Sun’s location within the Milky Way and the rotation speed at the Galactocentric distance of the Sun are irrelevant, because we are considering a local population and, thus, our simulation is scalable to more recent estimates (e.g., Gravity Collaboration et al., 2019).
@@
The simulated white dwarf spatial distribution follows double-decreasing exponential profiles in the Galactocentric cylindrical coordinates, and , with given scale length and height, and , for the disk components. We adopt 250 pc and 2.6 kpc for the thin disk, and 1.5 kpc and 3.5 kpc for the thick disk (Torres et al., 2019b). An isothermal distribution, with a local constant density, is assumed for the halo members.
@@
Regarding the velocity distribution, we adopted a Schwarzschild distribution where each Galactic component follows a Gaussian distribution that reproduces the average quantities observed for white dwarfs with respect to the heliocentric reference frame (Table LABEL:table:vel-gauss). That is, the adopted mean values are those corresponding to the observed sample following the classification provided by Torres et al. (2019b), with the astrometry updated to Gaia EDR3, while the assigned velocity dispersion of the thin disk stars is a function of their age, for which we adopted the average values of the full sample studied by Raddi et al. (2022) (see their table 8). As the adopted values are drawn from observed samples, they naturally account for the asymmetric drift that is typical of each given sub-population. We note that our synthetic population accounts for the vertex deviation, i.e. the correlation between and velocity components. The covariance of the -components for the observed stars are listed in Table LABEL:table:vel-gauss. From the observed sample, we measure small correlations (, , ) that correspond to a deviation of less than in the UV plane (positive for the thin disk, negative for the thich disk and halo). A larger value, of around , was found by Rowell & Kilic (2019) that, however, selected a white dwarf sample extending up to 250 pc. Moreover, when modeling the kinematic of the 100-pc sample, we did not account for purely local peculiar velocity groups such as the Hercules stream, which was shown to cause a strong overdensity of white dwarfs belonging to the thick disk in the -plane that are lagging behind the LSR (Torres et al., 2019a). By not accounting for the Hercules stream, it is expected that our simulated thick disk population will have a slower average velocity with respect to the observed population.
@@ 4 Orbit integration
(Bovy & Rix, 2013; Bovy, 2015), for orbit integration of both the observed and simulated populations. This is a static, axisymmetric potential, which does not include perturbative effects due to the presence of the Galactic bar and spiral arms. It consists of three time independent components, whose relative amplitudes are scaled to match observed properties of the Milky Way (Bovy, 2015): a spherical bulge modeled as a power-law density profile with an exponential cutoff (Bovy, 2015), an axisymmetric disk (Miyamoto & Nagai, 1975), and a dark-matter halo (Navarro et al., 1997). We note that the bulge has a lower mass with respect to other works (Bovy & Rix, 2013), implying reduced gravitational forces for objects that eventual approach the Galactic center. Further details on the scaling of the relative Galactic components are presented in Bovy (2015).
Orbits have been integrated in a time lapse of 2 Gyr with time-steps. This period corresponds to almost 9 complete orbits of the Sun around the Galactic center (about 224 Myr), ensuring the statistical significance of our analysis. Furthermore, in the analysis of Solar neighborhood’s lifetime (Section 5.1.1), we performed a shorter integration for 15 Myr with 1 000 time-steps.
@@ 5 Results
As the Sun orbits the Galactic center, the Solar neighborhood also moves. In other words, we are adopting a Lagrangian definition for our neighborhood space. It is expected, therefore, to have a flux of objects entering and exiting this volume. We emphasize that the results discussed in this section are valid on timescales shorter than the Milky Way’s lifetime and describe the average properties of a white dwarf sample in the Solar neighborhood.

### v2

# Kinematic origin of white dwarfs in the solar neighborhood

@@ Abstract
Context. White dwarfs are considered to be efficient cosmochronometers. Thanks to the recent space-borne mission Gaia, a nearly complete sample up to about 100 parsecs from the Sun has been compiled. However, the Galaxy, as a dynamic system, implies that these objects may have very diverse origins. It is therefore of paramount importance to characterize the origins of white dwarfs from the different Galactic structure components found in our solar neighborhood.
Aims. We aim to compute the Galactic orbits for white dwarfs of the thin and thick disk, as well as the halo components observed in our solar neighborhood. On the basis of these determinations, we analyze the most probable regions of the Galaxy where they could have formed, along with the distribution of their orbital parameters and the observational biases introduced when constructing the local sample.
Methods. We used a detailed Galactic orbit integration package, in conjunction with a detailed population synthesis code specifically designed to replicate the different Galactic components of the white dwarf population. Synthetic stars were generated based on the current observational sample and their orbital integration allowed for the reconstruction of the population’s history.
Results. Our kinematic analysis of the white dwarf population reveals the ephemeral nature of the concept of the solar neighborhood, as the majority of thin-disk, thick-disk, and halo white dwarfs will have left our 100 pc neighborhood in approximately 3.30 Myr, 1.05 Myr, and 0.6 Myr, respectively. Moreover, the spatial distribution of the integrated thin-disk orbits suggests that 68% of these stars were formed at less than 1 kpc from the Sun, while most of the thick-disk members have undergone radial disk migration. Halo members are those stars that typically belong to the “inner halo,” given that their orbits mostly planar and do not extend beyond –25 kpc. Despite the observational bias, which mostly affects the oldest stars in the thick disk and halo, we show that the wider distribution of orbital parameters is well represented by the sample.
Conclusions. The solar neighborhood is a transitory concept, whereby its current population of white dwarfs originates from larger regions of the Galaxy. This fact must be taken into account when analyzing the overall properties of such a population, such as its age distribution, metallicities, and formation history. Even so, the kinematic properties observed by recent missions such as Gaia are representative of the total population up to a distance of approximately 500 pc.
@@
The high-precision astrometry provided by the European Space Agency’s Gaia mission (Gaia Collaboration et al., 2016) has opened up a wide parameter space to Galactic studies, enabling the mapping of stellar kinematics over a large volume of the Milky Way, extending up to 6 kpc towards the inner and outer Galactic radial directions (Gaia Collaboration et al., 2018). The complementary information that is contributed by precise measurements of chemical abundances for ever-growing samples of stars is helping to shape the correlations between kinematics and stellar chemistry that are key to unraveling the past history of our Galaxy (Gaia Collaboration et al., 2023). The improved astrometry delivered by the Gaia Early Data Release 3 (EDR3; Gaia Collaboration et al., 2021) has confirmed the interplay between the kinematics of local structures, such as moving groups (Antoja et al., 2008), as well as that of more distant ones (Ramos et al., 2018; Khanna et al., 2019) in the Galactic anti-center direction and away from the mid-plane of the Galactic disk. This implies a connection with the recent passing of the Sagittarius dwarf galaxy (Ruiz-Lara et al., 2020) and interactions with the Large Magellanic Cloud (Antoja et al., 2018; Laporte et al., 2018).
@@
Pioneering works focused on the kinematic analyses of white dwarf samples in the pre-Gaia era (e.g., Silvestri et al., 2001; Pauli et al., 2006; Wegg & Phinney, 2012; Sion et al., 2014; Anguiano et al., 2017). However, more recently, by means of accurate Gaia parallaxes and proper motions, it has been possible to associate white dwarfs to known stellar streams (Torres et al., 2019a), to measure the reflex solar motion (Rowell & Kilic, 2019), to assign their membership to Galactic components within 100 pc (Torres et al., 2019b), and to accurately measure their age-velocity dispersion relation (Cheng et al., 2019; Raddi et al., 2022).
In this work, we model the kinematics of white dwarfs with a synthetic population by integrating their Galactic orbits in a static potential and using the present-day boundary conditions that have been obtained from the Gaia data. While a static potential is far from being a realistic approximation, the kinematics of local stars carries the imprint of resonances and radial mixing due to the Galactic structure and dynamics (Sellwood & Binney, 2002; Dehnen, 2000; Hunt et al., 2019), thus enabling a statistical study of the distributions that characterize their orbital parameters. The crossing-time of white dwarfs that are passing through the solar neighborhood is numerically quantified, as it is known to introduce a bias, although small, in the average properties of classes of local stars (Mayor et al., 1977). Moreover, we analyze the spatial distributions obtained by integrating for several orbits around the Galaxy as a statistical method for linking the observed white dwarf populations to the birth sites of their main sequence progenitors.
In the following sections, we introduce the properties of the observed and simulated samples (Sects. 2 and 3, respectively), describe the orbit integration framework (Sect. 4) and present a detailed analysis of the results (Sect. 5). In the last section, we conclude with a summary and the expectations for future developments.
@@ 2 Observed sample
The analysis of the white dwarf population provided by Gaia reveals that the largest sample that can be considered as effectively complete is the 100 pc volume-limited sample (Jiménez-Esteban et al., 2018). Thus, we started from the sample analyzed in Torres et al. (2019b), in which a random forest (RF) algorithm was employed to classify the white dwarf population within 100 pc of the Sun into its Galactic components. The sample was selected by applying color-magnitude cuts to the Hertzsprung–Russell diagram and adding specific selection criteria, such as photometric and astrometric relative errors under 10 per cent (see Sect. 4 from Torres et al., 2019b, for a detailed description). Then we cross-matched it with the Gaia EDR3, providing updated astrometry for the same sample.
@@
It is worth noting here that in order to eliminate possible contaminants in high-crowded areas, the observed sample analyzed in Torres et al. (2019b) did not include objects with proper motions corresponding to projected tangential velocities below km s-1. This fact should be taken into account in our analysis because it may introduce a slight bias for low-speed objects that are absent from the observed sample.
@@
The observed sample was thus classified using a RF algorithm applied to a 8D space (equatorial coordinates, parallax, proper motion components, and photometric magnitudes). The updated sample, with Gaia EDR3 astrometry, delivers 11,568 thin-disk, 1317 thick-disk, and 89 halo white dwarf candidates, implying a reduction of 6 % for each of the populations with respect to the number of candidates reported by Torres et al. (2019b). Such a small difference is attributed to stars that are now outside the 100 pc volume. We note that we did not search for new 100-pc candidates in EDR3. Although radial velocity measurements are absent in the Torres et al. (2019b) sample, it may be considered the largest, nearly complete, sample of white dwarfs, thus justifying a reliable analysis of the kinematics of the Galactic components of this population.
@@
Since most of the known white dwarfs do not have accurate radial velocity measurements, with very few of them having suitable high-resolution observations (Napiwotzki et al., 2020), a radial velocity of = 0 km s-1has been usually adopted in kinematic studies of this population. Alternative methods to overcome this issue have been proposed (Dehnen & Binney, 1998); however, they are not exempt from certain limitations (McMillan & Binney, 2009). In order to see what is the effect of considering the radial velocity zero, Torres et al. (2019b), analyzed the velocity component distributions in the Toomre diagram for the synthetic sample of the different Galactic components by subtracting the . Their analysis showed an average bias of 10 km s-1 and up to 74 km s-1 for the average component of the disk and halo stars, respectively, when the km s-1. The bias resulting by the assumption of zero radial velocity for white dwarfs is estimated to cause a reduction by a factor of one-third in the three velocity components of the heliocentric Cartesian reference system (e.g., Rowell & Kilic, 2019). Although it is beyond the scope of the present work to retrieve the full 3D velocity space of the white dwarf population, we address this issue in Section 5.2.2, where we attempt to correct for this bias using our synthetic model.
3 Synthetic white dwarf population
@@
Halo | ||||
Notes: T∗ is the total age of the star, which is the sum of the cooling time | ||||
of the white dwarf and the age of the progenitor. The velocities are given | ||||
in the heliocenric reference system. |
@@
In addition to the observational sample described above, we produced a synthetic sample of white dwarfs to address potential observational biases or deficiencies in the observational sample. We took advantage of the consolidated population synthesis code based on Monte Carlo (MC) techniques described in Torres et al. (2019b), which has been widely used over the last few decades in the study of the disk, halo, Galactic bulge, open clusters, and globular clusters (e.g., García-Berro et al., 1999; Torres et al., 2002; García-Berro et al., 2010; Cojocaru et al., 2015; Torres et al., 2015; Torres & García-Berro, 2016; Torres et al., 2018)
@@
The MC algorithm provides the initial position of each star within a 100 pc radius around the Sun and its age. In the following, we provide some of the basic ingredients used in our population synthesis code. According to a certain star formation rate (SFR), a probabilistic distribution determines the the ages of stars belonging to each Galactic component. We assumed a uniform SFR for the thin disk that lasts for a lifetime of , along with Gaussian distributions for the thick disk (, 1 Gyr) and the halo (, 0.5 Gyr), with 9.2, 10.0 and 12.5 Gyr, respectively (Torres et al., 2019b).
@@
The local standard of rest (LSR) is the chosen reference frame to describe the positions and velocities of the white dwarfs. It is instantaneously located on the Sun at kpc from the Galactic center, performing a circular orbit rotating at km s-1, which is nearly the average rotation velocity of stars in the solar neighborhood (Bovy et al., 2012). We placed the Sun at pc above the Galactic plane (e.g., Jurić et al., 2008). Nevertheless, the actual values for the Sun’s location within the Milky Way and the rotation speed at the Galactocentric distance of the Sun are irrelevant because we are considering a local population, thus, our simulation is scalable to more recent estimates (e.g., Gravity Collaboration et al., 2019).
@@
The simulated white dwarf spatial distribution follows double-decreasing exponential profiles in the Galactocentric cylindrical coordinates, and , with given scale length and height, and , for the disk components. We adopted 250 pc and 2.6 kpc for the thin disk and 1.5 kpc and 3.5 kpc for the thick disk (Torres et al., 2019b). An isothermal distribution, with a local constant density, was assumed for the halo members.
@@
Regarding the velocity distribution, we adopted a Schwarzschild distribution where each Galactic component follows a Gaussian distribution that reproduces the average quantities observed for white dwarfs with respect to the heliocentric reference frame (Table LABEL:table:vel-gauss). That is to say, the adopted mean values are those corresponding to the observed sample following the classification provided by Torres et al. (2019b), with the astrometry updated to Gaia EDR3, while the assigned velocity dispersion of the thin-disk stars is a function of their age, for which we adopted the average values of the full sample studied by Raddi et al. (2022) (see their Table 8). As the adopted values were drawn from observed samples, they naturally account for the asymmetric drift that is typical of each given sub-population. We note that our synthetic population accounts for the vertex deviation, namely, the correlation between and velocity components. The covariance of the -components for the observed stars are listed in Table LABEL:table:vel-gauss. From the observed sample, we measured small correlations (, , ), which correspond to a deviation of less than in the UV plane (positive for the thin disk, negative for the thich disk and halo). A larger value of around was found by Rowell & Kilic (2019); however, this value is based on a white dwarf sample extending up to 250 pc. Moreover, when modeling the kinematic of the 100-pc sample, we did not account for purely local peculiar velocity groups, such as the Hercules stream, which was shown to cause a strong overdensity of white dwarfs belonging to the thick disk in the -plane that are lagging behind the LSR (Torres et al., 2019a). By not accounting for the Hercules stream, it is expected that our simulated thick disk population will have a slower average velocity with respect to the observed population.
@@ 4 Orbit integration
(Bovy & Rix, 2013; Bovy, 2015) for the orbit integration of both the observed and simulated populations. This is a static, axisymmetric potential, which does not include perturbative effects due to the presence of the Galactic bar and spiral arms. It consists of three time-independent components, whose relative amplitudes are scaled to match observed properties of the Milky Way (Bovy, 2015): a spherical bulge modeled as a power-law density profile with an exponential cutoff (Bovy, 2015), an axisymmetric disk (Miyamoto & Nagai, 1975), and a dark-matter halo (Navarro et al., 1997). We note that the bulge has a lower mass with respect to values given in other works (Bovy & Rix, 2013), implying reduced gravitational forces for objects that eventual approach the Galactic center. Further details on the scaling of the relative Galactic components are presented in Bovy (2015).
Orbits were integrated in a time lapse of 2 Gyr with time-steps. This period corresponds to almost nine complete orbits of the Sun around the Galactic center (about 224 Myr), ensuring the statistical significance of our analysis. Furthermore, in the analysis of the solar neighborhood’s lifetime (Section 5.1.1), we performed a shorter integration for 15 Myr with 1 000 time-steps.
@@ 5 Results
As the Sun orbits the Galactic center, the solar neighborhood also moves. In other words, we have adopted a Lagrangian definition for our neighborhood space. Therefore, we expect to have a flux of objects entering and exiting this volume. We emphasize that the results discussed in this section are valid on timescales shorter than the Milky Way’s lifetime; hence, they describe the average properties of a white dwarf sample in the solar neighborhood.